OpenStudy (anonymous):

word problem: Anna is covering the circular pool with a heavy-duty coer for the winter. The pool has a diameter of 24 feet. The cover extends 12 inches beyond the edge of the pool, and a rope runs along the edge of the cover to secure the cover in place. What is the area of the cover? What is the length of the rope?! plz help!

OpenStudy (campbell_st):

the diameter of the cover will be 26 ft so the radius is r = 13  the area will be \[A = \pi \times13^2\] the rope length wil be \[C = \pi \times 26\]